OpenStudy (anonymous):

What is the limit of 3(1-tan x)sec x as it goes to pi/4?? Isn't it 2.94

OpenStudy (anonymous):

I plugged in pi over 3 into the equation, I get 2.95, and it says it's wrong.

OpenStudy (anonymous):

how do you get anything but 0?

OpenStudy (anonymous):

http://s11.postimage.org/itf544cfl/Problem2.png

OpenStudy (anonymous):

tan(pi/4) =1

OpenStudy (anonymous):

ahhh, I accidentally had the mode on my calculator set to 'degree' instead of 'radians'

OpenStudy (anonymous):

but even with radians, it gives me -1.24 as the final answer.

OpenStudy (anonymous):

Which is wrong....

OpenStudy (anonymous):

I have one more attempt on this problem >.>, otherwise it counts that I got it wrong.

OpenStudy (anonymous):

"x as it goes to pi/4??"

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

x-> pi/4 ?

OpenStudy (anonymous):

yeah, http://s11.postimage.org/itf544cfl/Problem2.png thats the problem

OpenStudy (anonymous):

There is not a need for L'hopitals rule considering that it's not something like infinity times infinity etc. The answer should be -1.24..

OpenStudy (anonymous):

Am I wrong?

OpenStudy (anonymous):

don't know what the problem is, so it's impossible to say.

OpenStudy (anonymous):

Is the answer I'm getting correct? -1.24?

OpenStudy (anonymous):

The website could be messing up.

OpenStudy (anonymous):

A few students had that as well.

OpenStudy (anonymous):

not clicking your hilarious link.

OpenStudy (anonymous):

?? That's a screenshot of the problem..

OpenStudy (anonymous):

'Hilarious' link? lol?

OpenStudy (anonymous):

cool story. upload it to OS.

OpenStudy (anonymous):

OS?

OpenStudy (anonymous):

Oh to openstudy

OpenStudy (anonymous):

there.

OpenStudy (anonymous):

http://i.imgur.com/CMgMz.gif

OpenStudy (anonymous):

\[\huge \huge \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \]

OpenStudy (anonymous):

Limit is zero

OpenStudy (anonymous):

how did you get 0?

OpenStudy (anonymous):

ah I think I got it. 1 - tan ( pi / 4 ) is 1- 1..so anything times 0..= 0..thats y?

OpenStudy (anonymous):

\[ \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 }\frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 }\frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 }\frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 } \frac {\pi }{4 }\]

OpenStudy (anonymous):

Ignore the 3 and sec x. lim 1-tanx as x approaches pi/4 equals zero because tan of pi/4 equals one

OpenStudy (anonymous):

Ah, I see. Very weird, the calculator kept giving me -1.24 o.O

OpenStudy (anonymous):

Okay, 0 worked :). Thanks